OpenStudy (anonymous):

Factor the following trinomial expression by a method of your choice: 3x2 + 7x-10. A. (3x + 2)(x - 5) B. (3x + 5)(x - 2) C. (3x + 1)(x - 10) D. (3x + 10)(x - 1)

OpenStudy (aripotta):

multiply the first and third terms. what do you get?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

Trying to factor by splitting the middle term 2.1 Factoring 3x2+7x-10 The first term is, 3x2 its coefficient is 3 . The middle term is, +7x its coefficient is 7 . The last term, "the constant", is -10 Step-1 : Multiply the coefficient of the first term by the constant 3 • -10 = -30 Step-2 : Find two factors of -30 whose sum equals the coefficient of the middle term, which is 7 . -30 + 1 = -29 -15 + 2 = -13 -10 + 3 = -7 -6 + 5 = -1 -5 + 6 = 1 -3 + 10 = 7 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -3 and 10 3x2 - 3x + 10x - 10 Step-4 : Add up the first 2 terms, pulling out like factors : 3x • (x-1) Add up the last 2 terms, pulling out common factors : 10 • (x-1) Step-5 : Add up the four terms of step 4 : (3x+10) • (x-1) Which is the desired factorization Final result : (x - 1) • (3x + 10)

OpenStudy (aripotta):

10 is negative, so you get -30. now you wanna find two numbers that when multiplied together get -30, and when added together get 7, the middle term.

OpenStudy (anonymous):

D @Aureyliant

OpenStudy (aripotta):

...

OpenStudy (anonymous):

Correct! @cassiesilva0628 please remember to ask if you don't understand something... you won't always have access to openstudy on tests and/or other things school related.

OpenStudy (anonymous):

@AriPotta Thank you. :)

OpenStudy (anonymous):

okay thank you so much for all you help @Aureyliant

OpenStudy (anonymous):

No problem, but @AriPotta deserves the medal more than I do because he/she was answering the Q first, and was actually engaging you rather than I was. :)